OpenStudy (anonymous):

if a=b 2a.a-2a.b=ab-2ab+a.a 2(a.a-a.b)=a.a-ab 2=1 ??!!!

OpenStudy (anonymous):

How is that ? 2=1 ?? :D

OpenStudy (anonymous):

I delete a.a-a.b from both I divised

OpenStudy (anonymous):

2(aa-ab)/(aa-ab)=1(aa-ab)/(aa-ab) 2=1 !!!

OpenStudy (anonymous):

Good u get it :D

OpenStudy (anonymous):

oh i didn't know it was a quiz, sorry

OpenStudy (anonymous):

no at least u got it :D we can't divise on 0 :D

OpenStudy (anonymous):

2a.a-2a.b=ab-2ab+a.a 2a^2-2a^2=a^2-2a^2+a^2 0=0 2(a.a-a.b)=a.a-ab 2(a^2-a^2)=a^2-a^2 2(0) =0 0=0 2=1 ??!!!

OpenStudy (anonymous):

is that what you need?

OpenStudy (anonymous):

I knew its wrong lol :D